OpenStudy (anonymous):

Suppose a particular jet needs to attain a speed of 220 mph to take off. If it can accelerate from 0 to 220 mph in 45 seconds, how long must the runway be? Assume constant acceleration. Note: 1 mph = 22/15 ft/sec. (answer in feet)

OpenStudy (ash2326):

@Ldaniel are you here

OpenStudy (anonymous):

is 7260ft the right answer?

OpenStudy (anonymous):

The answer should be 7260 feet. Reason: s=ut+1/2*a*t^2 u=initial vel=0 s= distance a= acceleration t= time taken Therefore a= (v-u) / t = (220-0) /45 = 220/45 s= 0+ (1/2) * (220/45)*(45)^2 Simplifying, you get s= 4950 miles Converting into feet by multiplying by 22/15 ft/sec You will get s=7260 feet @Siddharth18 , you are right :)

OpenStudy (anonymous):

yes i agry with him